OpenStudy (xbeenchantedx):

Completely simplify the expression below. All denominators should be rationalized. 4+sqrt (7)/sqrt(20)

OpenStudy (xbeenchantedx):

\[\frac{ 4+\sqrt{7} }{ \sqrt{20} }\]

OpenStudy (luigi0210):

Multiply by \(\sqrt{20}\)

OpenStudy (xbeenchantedx):

I did and I ended up with \[\frac{ 4\sqrt{20} + \sqrt{140}}{ 20 }\]

OpenStudy (anonymous):

\[\frac{1}{10} \left(4 \sqrt{5}+\sqrt{35}\right) \]